# PSBD: Prediction Shift Uncertainty Unlocks Backdoor Detection

### v1

@@ 4.3 From Insight to Innovation: Prediction Shift Backdoor Detection
Even with dropout enabled, the predicted labels for some clean data remain unchanged before and after applying dropout, although their prediction confidence changes significantly, leading to the potential misclassification of this clean data as backdoor data. To quantify the change in prediction confidence rather than the change in labels as defined in Equation (1), we introduce a new and more fine-grained measure of predictive uncertainty, Prediction Shift Uncertainty (PSU). PSU computes the difference between the predicted class confidence without dropout and the average predicted class confidence across $k$ dropout inferences to quantify the intensity of PS:
$\begin{split}c=\mathop{\arg\max}_{c\in\mathcal{C}}{P}(\mathbf{x};{\boldsymbol{\theta}}),\ \phi_{PSU}(\mathbf{x})={P}_{c}(\mathbf{x};{\boldsymbol{\theta}})-\frac{1}{k}\sum\limits_{i=1}^{k}{P}_{c}(\mathbf{x};p,{\boldsymbol{\theta}_{i}^{{}^{\prime}}})\end{split}$ (2)
@@
Similar to pilot studies, the appropriate dropout rate $p$ is a crucial factor in the dropout-based uncertainty method and is challenging to determine without knowledge of backdoor attacks. A reasonable $p$ is selected when the PS of clean data (both clean training data and clean validation data) achieves a relatively strong intensity, and that of backdoor data remains relatively weak. However, due to a lack of backdoor knowledge, we cannot directly compute the PS of backdoor data. Therefore, based on the definition of $\sigma$ provided in Equation (1), we propose an adaptive selection strategy for $p$. Specifically, we identify the $p$ where the $\sigma$ of clean validation data approach to a high value (0.8 in our experiments), while the difference between the $\sigma$ of the entire training data and that of the clean validation data reaches its maximum.
@@
(a) w/o data augmentation
(b) w/ data augmentation
Figure 4: The shift ratio curves of BadNets without and with data augmentation in Tiny ImageNet.
As we mentioned above, clean data always shift from the origin prediction class to another specific class, while backdoor data often remain static. Consequently, the PSU of clean training data and clean validation data will be close and high, whereas the PSU of backdoor data will be small under an appropriate $p$. For suspicious data $\mathbf{x}$, it can be determined as malicious based on a defender-specified threshold $T$. If $PSU(\mathbf{x})<T$, it is classified as a backdoor sample. We set $T$ based on the close proximity of PSU values between clean training data and extra clean validation data. In other words, in the absence of knowledge regarding the backdoor attack, $T$ can be roughly regarded as the tolerable loss rate for clean training data. In all our experiments, $T$ is set to the 25th percentile PSU value of the whole clean validation data. Furthermore, we found that using data augmentation in model training significantly outperforms the non-augmented training approach on Tiny ImageNet. As shown in Figure 4, we observe that with data augmentation under Tiny ImageNet, there is a significant decrease of $\sigma$ for backdoor data. It indicates that the use of data augmentation can intensify neuron bias, especially when the model has a lack of generalization ability to recognize the more sophisticated features. Hence, we incorporate data augmentation during model training when the model’s generality is lacking. The specific workflow of the prediction shift backdoor detection (PSBD) method is given as follows:
  • First, we train the model using a standard supervised learning algorithm on the suspicious training dataset, employing common data augmentation techniques when the model lacks generalization ability (in our experiments, we select the most commonly used random flipping and random cropping).
  • Next, we select the dropout rate $p$ based on the adaptive selection strategy. Then, we select a late-stage model to calculate the PSU values for the suspicious training data and clean validation data, due to its enhanced data fitting capability and robust neuron bias paths. We use the model from the 95th epoch. In fact, we find that any model from the 90th epoch onward performs equally well.
  • Finally, $T$ is set to the 25th percentile value of the PSU of clean validation data. For suspicious data $\mathbf{x}$, we can determine it is malicious based on defender-specified threshold $T$. If $PSU(\mathbf{x})<T$, we view it as a backdoor sample.
@@ 5.1 Experiment Settings
We conduct all experiments on the CIFAR-10 [34] and Tiny ImageNet [36] datasets using the ResNet-18 [35] architecture (Note that a defender is free to choose any model). We randomly select 5% of the total quantity of whole poisoned training dataset from the original test sets as our extra clean validation data. Further details can be found in the Appendix C.1.
@@
We evaluate our PSBD method against seven representative backdoor attacks, namely BadNets [3], Blend [23], TrojanNN[22], Label-Consistent [24], WaNet [5], ISSBA [37] and Adaptive-Blend [38]. Data augmentation during the model training was employed exclusively for Adaptive-Blend on CIFAR-10 and all experiments on Tiny ImageNet to achieve an attack success rate exceeding 85%. Without sacrificing generality, in our all experiments the target class $y_{t}$ of backdoor data is class 0. More detailed settings are presented in Appendix C.2.
@@
We compare our method with five classic and state-of-the-art backdoor data detection methods, namely Spectral Signature [39], Strip[19], Spectre[40], SCAN [41], and SCP [18]. All five detection methods were implemented and evaluated on the CIFAR-10 dataset. For the Tiny ImageNet dataset, SCAN was excluded due to its computationally intensive matrix eigenvalue computations, which significantly increased processing time. We run three trials for each experiment of all methods and report the average results across all cases as the final result. Please refer to the Appendix C.4 for the implementation details.
Table 1: The performance (TPR(SD)/FPR(SD)) on CIFAR-10 and Tiny ImageNet. We mark the best result in boldface while the value with underline denotes the second-best. The failed cases (i.e., TPR < 0.8) are marked in gray. Adaptive-Blend attack has a 1% poisoning ratio on CIFAR-10 and 2% poisoning ratio on Tiny ImageNet, while other attacks have a 10% poisoning ratio.
@@
                                                            CIFAR-10        (All standard deviation values should be multiplied by $10^{-2}$)
Badnet 1.000(0.0)/0.104(0.3) 0.389(2.7)/0.512(0.3) 1.000(0.0)/0.113(0.6) 0.953(0.0)/0.450(0.0) 1.000(0.0)/0.009(1.5) 1.000(0.0)/0.205(0.5)
Blend 1.000(0.0)/0.135(2.5) 0.438(0.5)/0.507(0.1) 0.993(0.2)/0.118(0.4) 0.953(0.0)/0.450(0.0) 0.991(0.3)/0.000(0.0) 0.939(8.3)/0.244(6.8)
TrojanNN 0.983(1.6)/0.171(4.5) 0.302(20)/0.509(0.1) 0.996(0.4)/0.112(0.3) 0.950(0.0)/0.450(0.0) 1.000(0.0)/0.000(0.0) 0.921(7.8)/0.227(3.8)
Label-Consistent 0.992(1.4)/0.130(3.4) 0.447(0.3)/0.506(0.0) 0.994(0.2)/0.117(1.1) 0.953(0.0)/0.450(0.0) 0.979(0.2)/0.014(1.2) 0.889(5.7)/0.237(5.4)
WaNet 1.000(0.0)/0.116(1.8) 0.456(0.3)/0.505(0.0) 0.050(1.4)/0.101(0.5) 0.951(0.2)/0.450(0.0) 0.891(1.4)/0.034(2.2) 0.869(8.4)/0.251(4.7)
ISSBA 1.000(0.1)/0.113(1.9) 0.436(0.5)/0.507(0.1) 0.774(6.9)/0.120(0.7) 0.950(0.0)/0.450(0.0) 0.963(0.6)/0.011(1.9) 0.939(4.1)/0.290(2.9)
Adaptive-Blend 0.982(2.6)/0.184(3.6) 0.608(19)/0.145(0.2) 0.014(0.5)/0.069(0.4) 0.753(8.3)/0.144(0.1) 0.000(0.0)/0.023(4.0) 0.779(10)/0.246(5.9)
Average 0.994(0.8)/0.136(2.6) 0.439(6.2)/0.456(0.1) 0.689(1.4)/0.107(0.6) 0.923(1.2)/0.406(0.0) 0.832(0.4)/0.013(1.5) 0.905(6.3)/0.243(4.3)
                                                            Tiny ImageNet    (All standard deviation values should be multiplied by $10^{-2}$)
Badnet 0.989(1.0)/0.088(0.3) 0.480(0.1)/0.502(0.0) 0.841(3.8)/0.108(1.9) 0.522(0.0)/0.497(0.0) / 0.999(0.0)/0.271(0.5)
Blend 0.919(3.8)/0.108(0.7) 0.478(0.0)/0.502(0.0) 0.249(3.2)/0.086(1.1) 0.522(0.0)/0.496(0.1) / 0.551(1.2)/0.260(0.7)
TrojanNN 0.961(0.6)/0.222(0.1) 0.478(0.0)/0.502(0.0) 0.963(1.0)/0.104(1.0) 0.522(0.0)/0.497(0.0) / 0.972(1.1)/0.301(2.8)
Label-Consistent 0.839(5.0)/0.039(1.5) 0.478(0.0)/0.502(0.0) 0.460(16)/0.088(0.9) 0.522(0.0)/0.497(0.0) / 0.741(3.8)/0.187(1.0)
WaNet 0.959(2.5)/0.086(1.8) 0.478(0.0)/0.502(0.0) 0.087(3.4)/0.082(0.2) 0.522(0.0)/0.497(0.0) / 0.446(5.4)/0.254(1.0)
ISSBA 0.886(2.6)/0.209(1.0) 0.478(0.0)/0.502(0.0) 0.954(3.6)/0.097(0.5) 0.522(0.0)/0.497(0.0) / 0.691(3.8)/0.297(1.3)
Adaptive-Blend 0.949(2.9)/0.095(3.7) 0.392(0.1)/0.502(0.0) 0.210(16)/0.099(0.6) 0.621(0.0)/0.497(0.0) / 0.505(13)/0.218(1.2)
Average 0.929(2.6)/0.121(1.3) 0.466(0.0)/0.502(0.0) 0.538(6.6)/0.095(0.9) 0.536(0.0)/0.497(0.0) / 0.701(4.0)/0.255(1.2)
Metric.
To directly assess the effectiveness of detection methods, we employ common binary classification metrics: True Positive Rate (TPR) and False Positive Rate (FPR). Our evaluation prioritizes achieving a high TPR to ensure effective identification of backdoor samples, while simultaneously maintaining a low FPR to minimize erroneous deletion of clean samples. Values inside brackets represent standard deviations (SD).
@@ 5.2 Experiment Results
As shown in Table 1, PSBD demonstrates excellent backdoor detection performance across a wide range of attack scenarios, while effectively preserving a substantial amount of clean data. The results also demonstrate that PSBD achieves a substantial improvement in detection performance compared to the defense baselines. In contrast, all baseline methods fail in some evaluation attacks. Specifically, on the CIFAR-10 dataset, the Spectral Signature method failed to detect backdoor data under all attack scenarios, while simultaneously misclassifying a substantial amount of clean data as backdoor data. This suggests that when the trigger pattern is relatively large or complex, the spectral signature property may be difficult to capture. The Spectre method demonstrated relatively effective detection capabilities across most attack scenarios, but it also has a high FPR. This is not desirable in practice as it filters out a significant amount of clean training data, which can lead to issues like overfitting due to insufficient training data. Although the SCAN, Strip, and SCP methods exhibited relatively acceptable performance across most attack scenarios, achieving a relatively high TPR and a low FPR, their effectiveness deteriorated somewhat when confronted with attacks employing confusion strategy such as WaNet. This confusion strategy aims to disrupt the model by retaining an equal or greater proportion of confounding samples that contain the trigger pattern but are correctly labeled. This weakens the mapping strength between the trigger pattern and the target label. Building upon this strategy, the Adaptive-Blend attack employs a more intricate incomplete trigger, significantly diminishing the detection performance of all baseline methods. Furthermore, owing to the extensive matrix eigenvalue computations involved, SCAN suffered from substantial computational inefficiency.
@@
On the more challenging Tiny ImageNet, all baseline methods failed in most attack scenarios. This failure is likely due to the increased complexity of image features, which weakened their ability to capture the mapping between trigger pattern and target label. Encouragingly, our PSBD method maintained its effectiveness, ranking in the top two for all attacks except for TrojanNN (where it still achieved a TPR of 0.961). 5% poisoning ratio results can be found in Table A4. By employing dropout to diminish the prominence of most image features and utilizing robust neuron bias paths, PSBD effectively discerned the mapping from trigger pattern to target label.
@@
PSBD Is More Robust.
Despite the ostensible randomness introduced by dropout, our PSBD method paradoxically exhibits low variance due to the robustness of the PS phenomena and neuron bias effect for clean and backdoor data under the appropriate dropout $p$. In real-world scenarios, where defenders often lack specific information about the poisoned data and target labels used by attackers, the robustness of a defense mechanism is crucial.
@@ 6 Conclusion
In our study, we developed PSBD, a simple and effective method to detect backdoor attacks in neural networks by focusing on Prediction Shift under dropout conditions, leading to the concept of neuron bias. By analyzing changes in prediction confidence with dropout, PSBD effectively distinguishes between clean and backdoor data across multiple datasets and attack types. This research contributes a practical and effective solution to the challenge of backdoor attacks in DNNs, marking a notable advancement in the field of neural network security. Future efforts could explore extending the PSBD method to a broader range of domains, such as natural language processing or time-series analysis.
7 Acknowledgement
Ren Wang is supported by the NSF under Grants 2246157 and 2319243, and the ORAU Ralph E. Powe Junior Faculty Enhancement Award.

### v2

@@ 4.3 From Insight to Innovation: Prediction Shift Backdoor Detection
Even with dropout enabled, the predicted labels for some clean data remain unchanged before and after applying dropout, although their prediction confidence changes significantly. To quantify the change in prediction confidence rather than the change in labels as defined in Equation (1), we introduce a new and more fine-grained measure of predictive uncertainty, Prediction Shift Uncertainty (PSU). PSU computes the difference between the predicted class confidence without dropout and the average predicted class confidence across $k$ dropout inferences to quantify the intensity of PS:
$\begin{split}\phi_{PSU}(\mathbf{x})&={P}_{c}(\mathbf{x};{\boldsymbol{\theta}})-\frac{1}{k}\sum\limits_{i=1}^{k}{P}_{c}(\mathbf{x};p,{\boldsymbol{\theta}_{i}^{{}^{\prime}}}),\\ c&=\mathop{\arg\max}_{c\in\mathcal{C}}{P}(\mathbf{x};{\boldsymbol{\theta}})\end{split}$ (2)
@@
Similar to pilot studies, the optimal dropout rate $p$ is a crucial factor in the dropout-based uncertainty method and is challenging to determine without knowledge of backdoor attacks. A reasonable $p$ is selected when the PS of clean data achieves a relatively strong intensity, and that of backdoor data remains relatively weak. However, due to a lack of backdoor knowledge, we cannot directly compute the PS of backdoor data. Thus, based on the definition of $\sigma$ provided in Equation (1), we propose an adaptive selection strategy for $p$. Specifically, we identify the $p$ where the $\sigma$ of clean validation data approach to a high value (0.8 in our experiments), while the difference between the $\sigma$ of the entire training data and that of the clean validation data reaches its maximum.
@@
As we mentioned above, clean data always shift from the origin prediction class to another specific class, while backdoor data often remain static. Consequently, the PSU of clean training data and clean validation data will be close and high, whereas the PSU of backdoor data will be small under an appropriate $p$. For suspicious data $\mathbf{x}$, it can be determined as malicious based on a defender-specified threshold $T$. If $PSU(\mathbf{x})<T$, it is classified as a backdoor sample. We set $T$ based on the close proximity of PSU values between clean training data and extra clean validation data. In other words, in the absence of knowledge regarding the backdoor attack, $T$ can be roughly regarded as the tolerable loss rate for clean training data. In all our experiments, $T$ is set to the 25th percentile PSU value of the whole clean validation data. Furthermore, we found that using data augmentation in model training significantly outperforms the non-augmented training approach on Tiny ImageNet. It indicates that the use of data augmentation can intensify neuron bias, especially when the model has a lack of generalization ability to recognize the more sophisticated features. Hence, we incorporate data augmentation during model training when the model’s generality is lacking. The specific workflow of the prediction shift backdoor detection (PSBD) method is given as follows:
  • First, we train the model using a standard supervised learning algorithm on the suspicious training dataset, employing common data augmentation techniques when the model lacks generalization ability.
  • Next, we select the dropout rate $p$ based on the adaptive selection strategy. Then, we select a late-stage model to calculate the PSU values for the suspicious training data and clean validation data, due to its enhanced data fitting capability and robust neuron bias paths.
  • Finally, for suspicious data $\mathbf{x}$, we can determine it is malicious based on defender-specified threshold $T$. If $PSU(\mathbf{x})<T$, we view it as a backdoor sample. $T$ is set to the 25th percentile value of the PSU of clean validation data in all our experiments.
@@ 5.1 Experiment Settings
We conduct all experiments on the CIFAR-10 [23], GTSRB [40] and Tiny ImageNet [37] datasets using the ResNet-18 [17] architecture. Please note that a defender is free to choose any architecture, as the sole objective is to detect any potential backdoor data that may exist within the dataset. We randomly select 5% of the total quantity of whole poisoned training dataset from the original test sets as our extra clean validation data. Further details can be found in the Appendix D.1.
@@
We evaluate our PSBD method against seven representative backdoor attacks, namely BadNets [13], Blend [6], TrojanNN[28], Label-Consistent [44], WaNet [32], ISSBA [25] and Adaptive-Blend [35]. We examined two scenarios with poisoning ratios of 5% and 10%. The main paper discusses the 10% poisoning ratio in detail, while the results for the 5% poisoning ratio are presented in Appendix E. Without sacrificing generality, in all our experiments, the target class $y_{t}$ is set to class 0. Data augmentation during the model training was employed exclusively for Adaptive-Blend on CIFAR-10, GTSRB, and all experiments on Tiny ImageNet to achieve an attack success rate exceeding 85%. More detailed settings are presented in Appendix D.2. We also verify the robustness of PSBD against potential adaptive attacks in Appendix  F.
@@
We compare PSBD with five classic and state-of-the-art backdoor data detection methods, namely Spectral Signature(SS) [43], Strip [11], Spectre [16], SCAN [42], SCP [14] and CD-L [19]. All six methods were implemented and evaluated on the CIFAR-10 dataset. For GTSRB and Tiny ImageNet datasets, SCAN was excluded due to its computationally intensive matrix eigenvalue computations, which significantly increased processing time. We run 10 trials for each experiment of all methods and report the average results across all cases as the final result. We found that the variances are relatively small, so we ignored them. Please refer to the Appendix D.3 for the implementation details.
Metric.
To assess the effectiveness of detection methods, we employ common classification metrics: True Positive Rate (TPR) and False Positive Rate (FPR). Our evaluation prioritizes achieving a high TPR to ensure effective identification of backdoor samples, while simultaneously maintaining a low FPR to minimize erroneous deletion of clean samples. Values inside brackets represent standard deviations (SD). Moreover, the results for the area under receiver operating curve (AUROC) can be found in the Appendix G.
Table 1: The performance (TPR/FPR) on CIFAR-10, GTSRB and Tiny ImageNet. We mark the best result in boldface while the value with underline denotes the second-best. The failed cases (i.e., TPR < 0.8) are marked in gray. Adaptive-Blend attack has a 1%/1%/2% poisoning ratio on CIFAR-10/GTSRB/Tiny ImageNet, while other attacks have a 10% poisoning ratio. OOT indicates that the method did not finish within the allocated time limit.
@@
                 CIFAR-10
Badnet 1.000/0.104 0.389/0.512 1.000/0.113 0.953/0.450 1.000/0.009 1.000/0.205 0.998/0.158
Blend 1.000/0.135 0.438/0.507 0.993/0.118 0.953/0.450 0.991/0.000 0.939/0.244 0.976/0.156
TrojanNN 0.983/0.171 0.302/0.509 0.996/0.112 0.950/0.450 1.000/0.000 0.921/0.227 0.999/0.161
Label-Consistent 0.992/0.130 0.447/0.506 0.994/0.117 0.953/0.450 0.979/0.014 0.889/0.237 0.962/0.159
WaNet 1.000/0.116 0.456/0.505 0.050/0.101 0.951/0.450 0.891/0.034 0.869/0.251 0.863/0.144
ISSBA 1.000/0.113 0.436/0.507 0.774/0.120 0.950/0.450 0.963/0.011 0.939/0.290 0.965/0.157
Adaptive-Blend 0.982/0.184 0.608/0.145 0.014/0.069 0.753/0.144 0.000/0.023 0.721/0.257 0.432/0.167
Average 0.994/0.136 0.439/0.456 0.689/0.107 0.923/0.406 0.832/0.013 0.899/0.244 0.855/0.157
                 GTSRB
Badnet 0.987/0.202 0.476/0.502 0.999/0.096 0.524/0.497 OOT 1.000/0.344 0.911/0.193
Blend 0.910/0.207 0.476/0.502 0.897/0.093 0.524/0.497 OOT 0.286/0.337 0.462/0.199
TrojanNN 0.952/0.212 0.476/0.502 0.639/0.096 0.524/0.497 OOT 0.113/0.345 0.967/0.194
Label-Consistent 0.944/0.203 0.476/0.502 1.000/0.115 0.524/0.497 OOT 0.998/0.362 0.416/0.175
WaNet 0.996/0.115 0.476/0.502 0.037/0.109 0.524/0.497 OOT 0.129/0.306 0.031/0.182
ISSBA 0.999/0.211 0.476/0.502 0.725/0.092 0.524/0.497 OOT 0.584/0.339 0.705/0.197
Adaptive-Blend 0.899/0.194 0.299/0.392 0.004/0.094 0.750/0.388 OOT 0.071/0.332 0.028/0.158
Average 0.955/0.192 0.451/0.486 0.614/0.099 0.556/0.481 OOT 0.454/0.338 0.503/0.185
                 Tiny ImageNet
Badnet 0.989/0.088 0.480/0.502 0.841/0.108 0.522/0.497 OOT 0.999/0.271 0.462/0.176
Blend 0.919/0.108 0.478/0.502 0.249/0.086 0.522/0.496 OOT 0.551/0.260 0.874/0.175
TrojanNN 0.961/0.222 0.478/0.502 0.963/0.104 0.522/0.497 OOT 0.972/0.301 0.985/0.150
Label-Consistent 0.839/0.039 0.478/0.502 0.460/0.088 0.522/0.497 OOT 0.741/0.187 0.931/0.203
WaNet 0.959/0.086 0.478/0.502 0.087/0.082 0.522/0.497 OOT 0.446/0.254 0.577/0.151
ISSBA 0.886/0.209 0.478/0.502 0.954/0.097 0.522/0.497 OOT 0.691/0.297 0.978/0.137
Adaptive-Blend 0.949/0.095 0.392/0.502 0.210/0.099 0.621/0.497 OOT 0.651/0.190 0.331/0.176
Average 0.929/0.121 0.466/0.502 0.538/0.095 0.536/0.497 OOT 0.722/0.251 0.734/0.167
@@ 5.2 Experiment Results
As shown in Table 1, PSBD demonstrates excellent backdoor detection performance across a wide range of attack scenarios, while effectively preserving a substantial amount of clean data. The results also demonstrate that PSBD achieves a substantial improvement in detection performance compared to the defense baselines. In contrast, all baseline methods fail in some evaluation attacks.
@@
Specifically, on the CIFAR-10 dataset, the Spectral Signature method failed to detect backdoor data under all attack scenarios, while simultaneously misclassifying a substantial amount of clean data as backdoor data. This suggests that when the trigger pattern is relatively large or complex, the spectral signature property may be difficult to capture. The Spectre method demonstrated relatively effective detection capabilities across most attack scenarios, but it also has a high FPR. This is not desirable in practice as it filters out a significant amount of clean training data, which can lead to issues like overfitting due to insufficient training data. Although the SCAN, Strip, SCP, and CD-L methods exhibited relatively acceptable performance across most attack scenarios, achieving a relatively high TPR and a low FPR, their effectiveness deteriorated when confronted with attacks employing confusion strategy such as WaNet and Adaptive-Blend. This confusion strategy aims to disrupt the model by retaining an equal or greater proportion of confounding samples that contain the trigger pattern but are correctly labeled. On the GTSRB dataset, which contains a larger number of classes, the Spectral Signature and Spectre methods completely fail to detect backdoor data. The Strip, SCP, and CD-L methods exhibit good detection performance only in a few specific attack scenarios. In contrast, our PSBD method demonstrates strong detection capabilities across all attack scenarios, highlighting its robustness and generalizability.
On the more challenging Tiny ImageNet, all baseline methods failed in most attack scenarios. This failure is likely due to the increased complexity of image features, which weakened their ability to capture the mapping between trigger pattern and target label. Encouragingly, our PSBD method maintained its effectiveness, ranking in the top two for all attacks except for TrojanNN (where it still achieved a TPR of 0.961) and ISSBA. The slight performance degradation under ISSBA is likely due to the model’s insufficient ability to extract features from the data. This is reflected in the significantly lower clean accuracy of the model compared to that of the clean model. The clean accuracy of models can be found in the Appendix Table A3. By employing dropout to diminish prominent image features and utilizing robust neuron bias paths, PSBD effectively discerned the mapping from trigger pattern to target label.
@@ 6 Conclusion
In our study, we developed PSBD, a simple and effective method to detect backdoor samples in the training dataset by focusing on Prediction Shift phenomenon under dropout conditions, leading to the concept of neuron bias effect. By analyzing changes in prediction confidence with and without dropout, PSBD effectively distinguishes between clean and backdoor data across multiple datasets and attack types. This research contributes a practical and effective solution to the challenge of backdoor attacks in DNNs, marking a notable advancement in the field of neural network security. Future efforts could explore extending the PSBD method to a broader range of domains, such as natural language processing or time-series analysis.
Acknowledgement
This work is supported by the NSF under Grants 2246157 and 2319243. We are thankful for the computational resources made available through NSF ACCESS and Argonne Leadership Computing Facility.